# Adversarial Training: Enhancing Out-of-Distribution Generalization for Learning Wireless Resource Allocation

### v1

@@ Abstract
Deep neural networks (DNNs) have widespread applications for optimizing resource allocation. Yet, their performance is vulnerable to distribution shifts between training and test data, say channels. In this letter, we resort to adversarial training (AT) for enhancing out-of-distribution (OOD) generalizability of DNNs trained in unsupervised manner. We reformulate AT to capture the OOD degradation, and propose a one-step gradient ascent method for AT. The proposed method is validated by optimizing hybrid precoding. Simulation results showcase the enhanced OOD performance of multiple kinds of DNNs across various channel distributions, when only Rayleigh fading channels are used for training.
Index Terms:
Out-of-distribution generalization, adversarial training, unsupervised learning, hybrid precoding.
@@ I Introduction
Deep learning is widely used for optimizing resource allocation in wireless systems, by designing deep neural networks (DNNs) to learn the mapping from environment parameters (say channels) into optimized resources [1]. With closed-form expressions of objective function and constraints, DNNs can be trained in unsupervised manner, hence the computations for generating labels are avoided [2, 1].
Majority of existing works assume that training and test data are identically distributed. When the DNNs are tested on data with unseen distributions, their out-of-distribution (OOD) performance often degrades obviously. In wireless communications, the distribution of environment parameters, especially channels, may vary rapidly due to many factors such as the changed locations of users and scatterers [3]. Although DNNs can be fine-tuned online with transfer learning, meta-learning, etc., collecting and training with new datasets are time-consuming. Limited by the up-to-date data and computing resources, such adaptations are infeasible in highly dynamic wireless environments [3] where the channel statistics change in tens of seconds [4]. Hence, improving the OOD generalizability of DNNs is crucial for their practical use.
Only a few research efforts attempt to increase the OOD generalization performance of deep learning for wireless applications. A data generation method was proposed in [5], which however requires the knowledge of possible channel distributions for testing. A hyper-network was designed in [6] to generate weights for a deep unfolding network, thereby enhancing its OOD performance. It was reported in [7] that deep unfolding itself exhibits acceptable performance on unseen distributions. Nonetheless, deep unfolding needs task-specific design, which restricts their applicability.
@@
To achieve good OOD performance, distributionally robust optimization (DRO) [8] was introduced. In order to be well-generalized across a range of distributions, DRO optimizes the weights of a DNN for performing better on the worst-case distribution. Unfortunately, DRO is hard to be solved [8].
Adversarial training (AT) can be regarded as an easy-to-implement alternative to DRO [9, 10]. AT was originally proposed to defend against adversarial attacks. Specifically, AT algorithms, exemplified by [11], generate adversarial examples (i.e., the misleading samples with crafted perturbations) and adjust the weights of DNNs with these examples.
@@
Both DRO and AT have been investigated only in the context of supervised learning, and can not be applied to unsupervised learning. Defense for adversarial attacks has also been studied in supervised learning-based wireless solutions [12]. However, the viability of using AT to enhance OOD generalization remains unexplored in wireless communications.
@@
In this letter, we resort to AT for improving the OOD generalizability of the DNNs that are trained unsupervisedly for optimizing resource allocation. The major contributions are summarized as follows. (1) To find the worst-case distribution in unsupervised learning, we reformulate DRO and AT, enabling them to be aware of the deterioration of DNNs. (2) We propose a method, namely unsupervised AT (UAT), to solve the reformulated AT problem. (3) We evaluate the OOD performance gain from the UAT on diverse channel distributions by training several DNNs only on Rayleigh channels, for optimizing hybrid precoding that maximizes sum-rate (SR) under quality of service (QoS) constraints.
@@ II Preliminary
Denote ${\bf x}$ as the input and ${\bf y}^{*}$ as the label for supervised learning, where the pair $({\bf x},{\bf y}^{*})$ follows a given distribution $\mathcal{D}$ during training. A DNN $f_{\bf y}({\bf x};{\bm{\theta}})$ parameterized by ${\bm{\theta}}$ can produce an output ${\bf y}$. The accuracy of the output can be improved by optimizing ${\bm{\theta}}$ with a loss function $L({\bf y},{\bf y}^{*})$.
To obtain satisfactory OOD performance on unknown distributions, DRO, minimizing the expected loss on the worst-case distribution $\tilde{\mathcal{D}}$ that incurs the highest loss, is formulated as [8]
@@
Although solving (1) enables the DNN to achieve minimal loss over all distributions in $\mathbb{D}$, the impossibility of sampling from the unsolved $\tilde{\mathcal{D}}$ renders the problem intractable [8].
To defend against adversarial attacks, AT is formulated as follows to minimize the expected loss over the most “maliciously” perturbed inputs $\tilde{\bf x}$ [11],
@@
Differing from (1), the “optimization variable” in (2), $\tilde{\bf x}$, is a sample rather than a distribution, which is referred to as the adversarial example of ${\bf x}$. Since the expectation in (2) can be approximated by averaging over samples drawn from $\mathcal{D}$, this problem can be solved by alternately executing the following two steps [11]: (i) Find $\tilde{\bf x}$: It can be obtained using gradient ascent methods, guided by the partial derivative of loss function with respect to (w.r.t.) ${\bf x}$. (ii) Update ${\bm{\theta}}$: It can be optimized using back-propagation, with the loss averaged over all adversarial examples.
Previous studies have proved that the objective in (1) is upper bounded by the objective in (2) asymptotically (say Corollary 2 in [9], Theorems 1 and 2 in [10]). Consequently, DNNs trained by AT algorithms can also exhibit favorable OOD generalizability [10]. However, existing works never investigate DRO or AT for unsupervised deep learning.
@@ III Adversarial Training in Unsupervised Learning
In this section, we formulate DRO and AT for unsupervised learning, and propose a method for solving the AT problem.
@@ III-A Problem Reformulation
With environment parameter ${\bf x}$, resource allocation can be optimized from the following problem
@@
The optimal solutions for every value of ${\bf x}$ constitute a function ${\bf y}^{*}=f_{{\bf y}^{*}}({\bf x})$. When a DNN ${\bf y}=f_{\bf y}({\bf x};{\bm{\theta}})$ is trained to learn this function, ${\bf x}$ follows a distribution $\mathcal{D}$.
Unsupervised primal-dual learning (PDL) [1, 2] can be adopted to train the DNN, which can be formulated as
@@
where $\mathcal{L}({\bf x};{\bm{\theta}},{\bm{\xi}})$ is a Lagrangian function defined as follows
@@
and $f_{\lambda_{i}}({\bf x};{\bm{\xi}})$ is the $i$th element of an multiplier network $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$ [2].
To improve OOD performance, we consider DRO with a cost $F({\bf x},{\bf y})$,111“Cost” is a metric of interest. If one seeks to further guarantee the constraints on unseen distributions, $F({\bf x},{\bf y})+\sum_{i=1}^{N_{C}}\max(0,G_{i}({\bf x},{\bf y}))$ can be used as the cost, and the subsequent analysis is similar. which determines the worst-case distribution.
Analogous to (1), one can optimize the weights of DNNs on the distribution $\tilde{\mathcal{D}}$ with the highest cost, i.e., $\max_{\tilde{\mathcal{D}}\in\mathbb{D}}\mathbb{E}_{{\bf x}\sim\tilde{\mathcal{D}}}\left[F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))\right]$. However, $\tilde{\mathcal{D}}$ obtained in this way is NOT the worst-case distribution, because this objective cannot differentiate the cost increase caused by the higher optimal objective value and by the degraded generalization performance of $f_{\bf y}(\cdot\,;{\bm{\theta}})$ on the distributions in $\mathbb{D}$. The former is unrelated to generalizability. For example, for SR-maximization resource allocation, a decrease in channel strength inevitably increases the cost (i.e., reduces SR), but a DNN may not perform worse on low-strength channels. Optimizing the DNN on such channels may not be beneficial for OOD generalization.
To address this issue, [^2] This issue does not arise in supervised learning, where the cost is the same as the loss function, because the optimal cost/loss is always zero. we reformulate DRO as
@@
for training $f_{\bf y}(\cdot\,;{\bm{\theta}})$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})$ on the true worst-case distribution, where the DNN $f_{\bf y}(\cdot\,;{\bm{\theta}})$ has the largest gap in cost from the optimal solution instead of the highest cost.
Inspired by [10], we adopt AT as an realizable alternative to the DRO for enhancing OOD generalizability performance. By replacing $\tilde{\mathcal{D}}$ by $\tilde{\bf x}$, the AT problem for unsupervised PDL is formulated as
@@
III-B Unsupervised Adversarial Training (UAT) Method
To solve $P_{\rm AT}$, we first find the adversarial example $\tilde{\bf x}$ for every original sample ${\bf x}$ from (7b).
Recall the definitions of $\tilde{\bf x}={\bf x}+{\bm{\delta}}$ and $B_{\epsilon}({\bf x})$ in (2), (7b) can be transformed into a problem to optimize ${\bm{\delta}}$ as follows
@@ III-B Unsupervised Adversarial Training (UAT) Method
To find ${\bm{\delta}}$ (and hence find $\tilde{\bf x}$) with low complexity, we employ one-step gradient ascent method. By using the first-order Taylor polynomial, it can be seen that (8) has the same solution as the following problem when $\epsilon$ is very small,
@@
It is not hard to show that the optimal solution of (9) is
@@
To obtain the value of $\frac{\partial F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))}{\partial{\bf x}}-\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}({\bf x}))}{{\rm d}{\bf x}}$, we rewrite its first term by using the chain rule as
@@
where the values of partial derivatives $\frac{\partial F}{\partial{\bf x}}$ and $\frac{\partial F}{\partial{\bf y}}$ are computed with the value of ${\bf y}$ produced by $f_{\bf y}({\bf x};{\bm{\theta}})$, and $\frac{\partial f_{\bf y}({\bf x};{\bm{\theta}})}{\partial{\bf x}}$ can be obtained from the DNN due to its differentiability.
Obtaining the value of the second term is challenging owing to the unknown function $f_{{\bf y}^{*}}(\cdot)$. To cope with this difficulty, we invoke the envelope theorem [13]. Then, the gradient of the optimal objective can be obtained directly from (3) as
$\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}({\bf x}))}{{\rm d}{\bf x}}\!=\!\frac{\partial F}{\partial{\bf x}}+\lambda_{1}^{*}\!({\bf x})\frac{\partial G_{1}}{\partial{\bf x}}+\cdots+\lambda_{N_{C}}^{*}\!({\bf x})\frac{\partial G_{N_{C}}}{\partial{\bf x}},\!\!\!$ (12)
where $\lambda_{1}^{*}({\bf x}),\cdots,\lambda_{N_{C}}^{*}({\bf x})$ are the optimal Lagrange multipliers for the $N_{C}$ constraints in (3), which depend on ${\bf x}$.
Unlike (11), the values of partial derivatives in (12) are computed with the value of optimal solution ${\bf y}^{*}$. In practice, the value of ${\bf y}^{*}$ can be approximated by the output of ${\bf y}=f_{\bf y}({\bf x};{\bm{\theta}})$, since ${\bf x}$ herein is an in-distribution (ID) sample and $f_{\bf y}(\cdot;{\bm{\theta}})$ has been trained with all ID samples before UAT begins. Similarly, the values of $\lambda_{1}^{*}({\bf x}),\cdots,\lambda_{N_{C}}^{*}({\bf x})$ can be approximated by the outputs of multiplier network $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$.
@@
To solve $P_{\rm AT}$, the optimizations in (7b) and (7a) are alternately executed for $T$ iterations. In the first iteration, the adversarial examples are generated from the original samples using (14). In the subsequent iterations, the adversarial examples are generated from those obtained in the previous iteration. In this way, the range of perturbations can be expanded for facilitating OOD generalization, meanwhile in each iteration $\epsilon$ can be set as a small value for the Taylor approximation being accurate.
The proposed method is summarized in Algorithm 1, where Lines 8$\sim$15 are the steps specifically for UAT and other lines are for PDL. The lines marked with an asterisk (*) can be omitted if there are no constraints related to ${\bf x}$.
Remark 2: The training complexity of UAT is $T$ times higher than PDL. As a training method, UAT does not affect the inference complexity of DNNs.
@@
1:Input: number of epochs $E$, number of batches $B$, batch-size $N_{B}$, samples $\{{\bf x}_{n}\}_{n=1}^{BN_{B}}$, initial weight ${\bm{\theta}}$, learning rate $\eta_{1},\eta_{2}$, epoch to begin UAT $E_{A}$, number of adversarial iterations $T$, perturbation range $\epsilon$
2:* initialized weight of multiplier network ${\bm{\xi}}$
@@
7:         *$\mathcal{L}\!\leftarrow\!\mathcal{L}\!+\frac{1}{N_{B}}\!\!\sum_{n}\!\!\sum_{i=1}^{N_{C}}\!f_{\lambda_{i}}({\bf x}_{n};{\bm{\xi}})G_{i}({\bf x}_{n},f_{\bf y}({\bf x}_{n};{\bm{\theta}}))$
@@
9:         *${\bm{\xi}}\leftarrow{\bm{\xi}}+\eta_{2}\frac{\partial\mathcal{L}}{\partial{\bm{\xi}}}$
@@
13:                   $\tilde{\bf x}_{n}^{t}=\tilde{\bf x}_{n}^{t-1}\!+\epsilon\cdot d(\tilde{\bf x}_{n}^{t-1}),\!\forall n\!\in\!\mathcal{N}$triangleright see (14)
@@
15:                   *$\!\mathcal{L}\!\!\leftarrow\!\!\mathcal{L}\!+\!\frac{1}{N_{B}}\!\!\sum_{n}\!\!\sum_{i=\!1}^{\!N_{C}}\!\!f_{\!\lambda_{i}}\!(\!\tilde{\bf x}_{n}^{t};\!{\bm{\xi}}\!)G_{i}\!(\tilde{\bf x}_{n}^{t},\!f_{\bf y}(\tilde{\bf x}_{n}^{t};\!{\bm{\theta}})\!)$
16:                   ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{1}\frac{\partial\mathcal{L}}{\partial{\bm{\theta}}}$triangleright via back-propagation
17:                   *${\bm{\xi}}\leftarrow{\bm{\xi}}+\eta_{2}\frac{\partial\mathcal{L}}{\partial{\bm{\xi}}}$triangleright via back-propagation

### v2

@@ Abstract
[…]
@@ I Introduction
Deep learning is widely used for optimizing wireless resource allocation, by designing deep neural networks (DNNs) to learn the mapping from environment parameters (say channels) into optimized variables [1]. For wireless resource optimization with closed-form expressions of objective function and constraints, DNNs can be trained in an unsupervised manner. As a consequence, the labels obtained by solving computation-intensive numerical algorithms are no longer required, and constraint satisfaction can be controlled [1, 2].
The majority of existing works have assumed that training and test data are identically distributed. However, when the DNNs are tested on data drawn from unseen distributions, their out-of-distribution (OOD) performance often degrades obviously. In wireless communications, the distribution of environment parameters, especially channels, may vary rapidly due to many factors such as the changed locations of users and scatterers [3]. As a result, the distributions of training and test data may differ. A strong OOD generalizability enables DNNs to perform well across scenarios without further adaptation. This reduces the computational demands at wireless edge and aligns with the vision of green artificial intelligence (AI).
A few research efforts have attempted to enhance the OOD generalization of deep learning for wireless applications.
A data generation method was proposed in [4] to synthesize channel samples, based on an observation that DNNs trained on channels with proper power delay profiles are generalized well to other distributions when learning channel estimation. A data augmentation method was proposed in [5] after finding the major differences among channel distributions that are critical for channel feedback. These methods leverage the knowledge of certain tasks and are hard to be applied to other tasks.
It was reported that deep unfolding networks can achieve acceptable performance on unseen distributions when optimizing hybrid precoding[6] and power control[7], even without dedicated design for OOD generalizability. It was shown in [8] that the hyper-networks generating weights for the unfolding networks can achieve preferable OOD performance for signal detection when trained on multiple distributions. However, deep unfolding networks are problem- and algorithm-specific, which restricts their applicability.
Feature extractor (which is a DNN) was designed to capture the domain-invariant feature that does not change across distributions from each channel input, for learning reconfigurable intelligent surface configuration selection in [9] and indoor localization in [10]. The extracted feature was then fed into another DNN to produce the desired output. Nonetheless, the domain-invariant feature may not exist for all wireless problems, say precoding [11]. Meanwhile, samples from multiple distributions are required for training the feature extractor, which may not always be available in practice.
DNNs can be fine-tuned with transfer learning or meta-learning. Yet such adaptations are often infeasible in dynamic wireless environments [3] where the channel statistics change in tens of seconds [12]. This is because collecting and training with up-to-date samples are time-consuming and limited by the computing resource at wireless edge. In [13], a DNN for learning hybrid precoding was trained by meta-learning and evaluated for OOD generalization without adaptation, whose performance is only about half of that with sufficiently fine-tuning.
@@
To achieve good OOD performance, distributionally robust optimization (DRO) has been introduced into machine learning [14]. In order to be well-generalized across a range of distributions, DRO optimizes the trainable weights of a DNN to perform better on the worst-case distribution where the DNN deteriorates most severely. Unfortunately, DRO problem is hard to be solved since it involves the optimization of both the weights and the worst-case distribution [14].
Adversarial training (AT) methods, originally proposed to defend against adversarial attacks, have been used as an easy-to-implement alternative to DRO [15, 16]. Specifically, it was proved that the objective of DRO is upper bounded by the objective of AT asymptotically (say Corollary 2 in [15], Theorems 1 and 2 in [16]). Consequently, DNNs trained by AT can also exhibit favorable OOD generalizability [16].
@@
Both DRO and AT have been investigated only in the context of supervised learning, and can not be applied to unsupervised learning. The detection and defense for adversarial attacks has also been studied in supervised learning-based wireless AI [17]. However, the viability of using AT to enhance OOD generalization has not been explored for wireless communications.
@@
In this paper, we resort to AT for improving the OOD generalizability of the DNNs that are trained unsupervisedly for optimizing resource allocation. The major contributions are summarized as follows. (1) To find the worst-case distribution in unsupervised learning, we reformulate DRO and AT problems, enabling them to be aware of the deterioration of DNNs. (2) We propose a one-step gradient ascent algorithm, referred to as unsupervised AT (UAT), to solve the reformulated AT problem. (3) We evaluate the OOD generalization gain from the UAT on diverse channel distributions by training several DNNs only on a single channel distribution, for optimizing hybrid precoding that maximizes sum-rate (SR) under quality of service (QoS) constraints.
The proposed UAT has three advantages: (1) It does not require specific expert knowledge to generate or augment data. (2) It does not need to change the architectures of DNNs, in contrast to deep unfolding and domain-invariant feature extraction. (3) It does not rely on data diversity, eliminating the need to access or distinguish samples from different distributions, in contrast to meta-learning, training hyper-networks or feature extractors.
@@ II Preliminary
Denote $\bf x$} as the input and $\bf y$$^{*}$ as the label for supervised learning, where the pair $({\bf x},{\bf y}^{*})$ follows a given distribution (denoted by $\mathcal{D}$) during training. A DNN denoted by $f_{\bf y}({\bf x};{\bm{\theta}})$ can produce an output $\bf y$} for each $\bf x$}, where the trainable weight $\bm{\theta}$ is optimized through a loss function $L({\bf y},{\bf y}^{*})$ to encourage the output $\bf y$} to approach the label $\bf y$$^{*}$.
To obtain satisfactory OOD performance on unknown distributions, DRO minimizes the expected loss on the worst-case distribution (denoted by $\tilde{\mathcal{D}}$) that incurs the maximal loss, which is formulated as [14]
@@
Although solving the problem in (1) enables the DNN to achieve minimal loss over all distributions in $\mathbb{D}$, which improves its OOD generalizability, the impossibility of sampling from the unsolved $\tilde{\mathcal{D}}$ renders the problem intractable [14].
AT methods find adversarial examples and adjust the trainable weights of DNNs with these examples to prevent possible attacks. Specifically, AT problem minimizes the expected loss over the most maliciously perturbed inputs (denoted by $\tilde{\bf x}$) [18], which is formulated as follows
@@
Differing from (1), the “optimization variable” $\tilde{\bf x}$ in (2) is a sample rather than a distribution, which is referred to as the adversarial example of $\bf x$}. Since the expectation in (2) can be approximated by averaging over samples drawn from $\mathcal{D}$, this problem can be solved by alternating the following two steps [18]: (i) Find $\tilde{\bf x}$: It can be obtained using gradient ascent methods, guided by the partial derivative of loss function with respect to $\bf x$}. (ii) Update $\bm{\theta}$}: It can be optimized using back-propagation, with the loss averaged over all adversarial examples.
@@ III Adversarial Training in Unsupervised Learning
In this section, we formulate DRO and AT for unsupervised learning, and propose an algorithm for solving the AT problem.
@@ III-A Problem Reformulation
With environment parameter $\bf x$}, resource allocation $\bf y$ can be optimized from the following problem
@@
The optimal solutions for every value of $\bf x$} constitute a function ${\bf y}^{*}=f_{{\bf y}^{*}}({\bf x})$. When a DNN ${\bf y}=f_{\bf y}({\bf x};{\bm{\theta}})$ is trained to learn this function, $\bf x$} follows the distribution $\mathcal{D}$.
By introducing a multiplier network (denoted by $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$), primal-dual learning (PDL) [2, 1] can be adopted to train $f_{\bf y}(\cdot\,;{\bm{\theta}})$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})$ from the following problem
@@
where $\mathcal{L}({\bf x};{\bm{\theta}},{\bm{\xi}})$ is the Lagrangian function defined as
@@
and $f_{\lambda_{i}}({\bf x};{\bm{\xi}})$ represents the $i$th element of the output $\lambda_{1},\cdots,\lambda_{N_{C}}]^{T}=f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$.
After being well-trained, the resource allocation variable provided by $f_{\bf y}(\cdot\,;{\bm{\theta}})$ can achieve near-optimal objective value in (3) and satisfy the constraints with high probability [2, 1].
To improve OOD performance, we should introduce a cost function [^1] In supervised learning, the cost function is the same as the loss function, since the loss can indicate the performance of DNNs. to find the worst-case distribution for DRO, which measures how worse the DNN performs on a distribution. In this paper, we focus on pursuing better objective on unseen distributions, and thus select $F({\bf x},{\bf y})$ as the cost function. [^2] If one seeks to further guarantee the constraints on unseen distributions, $F({\bf x},{\bf y})+\sum_{i=1}^{N_{C}}\max(0,G_{i}({\bf x},{\bf y}))$ can be used as the cost function, and the subsequent analysis is similar.
Analogous to (1), one can optimize $\bm{\theta}$} on the distribution $\tilde{\mathcal{D}}$ with maximal cost, which is the solution of $\max_{\tilde{\mathcal{D}}\in\mathbb{D}}\mathbb{E}_{{\bf x}\sim\tilde{\mathcal{D}}}\left[F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))\right]$. However, $\tilde{\mathcal{D}}$ obtained in this way is NOT the worst-case distribution, because this objective cannot differentiate the cost increase caused by the higher optimal objective value and by the degraded generalization performance of $f_{\bf y}(\cdot\,;{\bm{\theta}})$ on the distributions in $\mathbb{D}$. The former is unrelated to generalizability. For example, for SR-maximization resource allocation, a decrease in channel strength inevitably increases the cost (i.e., reduces SR), but a DNN may not perform worse on low-strength channels. Optimizing the DNN on such channels is not beneficial for OOD generalization.
To address this issue, [^3] This issue does not arise in supervised learning, since the optimal value of the losses (such as cross-entropy, mean squared error) is always zero. we reformulate DRO for PDL as
@@
for training $f_{\bf y}(\cdot\,;{\bm{\theta}})$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})$ on the truely worst distribution, where the output of $f_{\bf y}(\cdot\,;{\bm{\theta}})$ yields the largest gap in cost from the optimal solution instead of the highest cost.
Inspired by the idea in [16], we adopt AT as a realizable alternative to the DRO for enhancing OOD generalizability performance. By replacing $\tilde{\mathcal{D}}$ by $\tilde{\bf x}$, the AT problem for PDL is reformulated as
@@
III-B Unsupervised Adversarial Training (UAT) Algorithm
To solve $P_{\rm AT}$, i.e., to train $f_{\bf y}(\cdot\,;{\bm{\theta}})$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})$ with found adversarial examples, (7b) and (7a) can be alternately optimized. To find accurate adversarial examples, both DNNs need to first be trained with in-distribution (ID) samples, namely, original samples provided in the training set.
In what follows, we show how to find the adversarial example $\tilde{\bf x}={\bf x}+{\bm{\delta}}$ for every original sample $\bf x$} from (7b).
From the definition of $B_{\epsilon}({\bf x})$ in (2), (7b) can be transformed into the following problem to optimize $\bm{\delta}$}
@@ III-B Unsupervised Adversarial Training (UAT) Algorithm
To find $\bm{\delta}$} (and hence find $\tilde{\bf x}$) with low complexity, we employ a one-step gradient ascent method. By using the first-order Taylor polynomial, it can be seen that (8) has the same solution as the following problem when $\epsilon$ is very small,
@@
The problem in (9) maximizes a linear objective over an Euclidean ball, whose optimal solution can be obtained as
@@
To obtain the value of the numerator in (10), we rewrite its first term by using the chain rule as
@@
where the values of partial derivatives $\frac{\partial F}{\partial{\bf x}}$ and $\frac{\partial F}{\partial{\bf y}}$ can be computed using the value of $\bf y$} produced by $f_{\bf y}({\bf x};{\bm{\theta}})$, and $\frac{\partial f_{\bf y}({\bf x};{\bm{\theta}})}{\partial{\bf x}}$ can be obtained from the DNN due to its differentiability.
Obtaining the value of the second term of the numerator is challenging, due to the unknown function $f_{{\bf y}^{*}}(\cdot)$. To cope with this difficulty, we invoke the envelope theorem [19], which provides the gradient of the optimal objective from (3), with no need for the closed-form expression of $f_{{\bf y}^{*}}(\cdot)$, as
$\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}({\bf x}))}{{\rm d}{\bf x}}=\frac{\partial F}{\partial{\bf x}}+\lambda_{1}^{*}\frac{\partial G_{1}}{\partial{\bf x}}+\cdots+\lambda_{N_{C}}^{*}\frac{\partial G_{N_{C}}}{\partial{\bf x}},$ (12)
where $\lambda_{1}^{*},\cdots,\lambda_{N_{C}}^{*}$ are the optimal Lagrange multipliers for the $N_{C}$ constraints in (3), which depend on $\bf x$}.
Unlike (11), the values of partial derivatives in (12) are computed with the value of optimal solution $\bf y$$^{*}$. In practice, the value of $\bf y$$^{*}$ can be approximated by $\bf y$} (i.e., the output of $f_{\bf y}({\bf x};{\bm{\theta}})$), since $\bf x$} herein is an ID sample and $f_{\bf y}(\cdot;{\bm{\theta}})$ has been trained with all ID samples before UAT begins. Similarly, the values of $\lambda_{1}^{*},\cdots,\lambda_{N_{C}}^{*}$ can be approximated by the outputs of multiplier network $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$.
@@
During UAT, (7b) and (7a) are alternately solved for $T$ iterations. In the first iteration, the adversarial examples are generated from the original samples using (14). In the subsequent iterations, the adversarial examples are generated from those obtained in the previous iteration. In this way, the range of perturbations can be expanded for facilitating OOD generalization, while in each iteration $\epsilon$ can be set as a small value for the Taylor approximation being accurate.
[…]
@@
1:Input: number of epochs $E$, number of batches $B$, batch-size $N_{B}$, samples $\{{\bf x}_{n}\}_{n=1}^{BN_{B}}$, initial value of weight $\bm{\theta}$}, learning rate $\eta_{1},\eta_{2}$, the epoch at which UAT begins $E_{A}$, number of adversarial iterations $T$, perturbation range $\epsilon$
2:* initial weight of multiplier network $\bm{\xi}$}
@@
7:   $\mathcal{L}\!\leftarrow\!\mathcal{L}\!+\frac{1}{N_{B}}\!\!\sum_{n}\!\!\sum_{i=1}^{N_{C}}\!f_{\lambda_{i}}({\bf x}_{n};{\bm{\xi}})G_{i}({\bf x}_{n},f_{\bf y}({\bf x}_{n};{\bm{\theta}}))$
@@
9:   ${\bm{\xi}}\leftarrow{\bm{\xi}}+\eta_{2}\frac{\partial\mathcal{L}}{\partial{\bm{\xi}}}$
@@
13:       $\tilde{\bf x}_{n}^{t}=\tilde{\bf x}_{n}^{t-1}\!+\epsilon\cdot d(\tilde{\bf x}_{n}^{t-1}),\!\forall n\!\in\!\mathcal{N}$ $\triangleright$ see (14)
@@
15:       $\mathcal{L}\!\!\leftarrow\!\!\mathcal{L}\!+\!\frac{1}{N_{B}}\!\!\sum_{n}\!\!\sum_{i=\!1}^{\!N_{C}}\!\!f_{\!\lambda_{i}}\!(\!\tilde{\bf x}_{n}^{t};\!{\bm{\xi}}\!)G_{i}\!(\tilde{\bf x}_{n}^{t},\!f_{\bf y}(\tilde{\bf x}_{n}^{t};\!{\bm{\theta}})\!)$
16:       ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{1}\frac{\partial\mathcal{L}}{\partial{\bm{\theta}}}$ $\triangleright$ via back-propagation
17:       ${\bm{\xi}}\leftarrow{\bm{\xi}}+\eta_{2}\frac{\partial\mathcal{L}}{\partial{\bm{\xi}}}$ $\triangleright$ via back-propagation